PA 12-2 Sarahs Organic Soap Company makes organic... Sarahs Organic Soap Company makes organic liquid soap. One of the raw
Answer #1

Daily demand d = 1500 kg

Annual demand D = 1500*5*52 = 390000

Ordering cost S = $56

Cost = $6 per kg

Holding cost = 15% = 0.15*4 = 0.9

a)

Optimal order quantity Q

2DS H

2 * 390000 56 0.9

Q = 6967 kg

b)

Order quantity Q = 3500 kg

Total Cost = Holding Cost + Ordering Cost = (Q/2)H + (D/Q)S

Total cost = (3500/2)*0.9 + (390000/3500)*56 = 7815

c)

Order quantity Q = 8500 kg

Total Cost = Holding Cost + Ordering Cost = (Q/2)H + (D/Q)S

Total cost = (8500/2)*0.9 + (390000/8500)*56 = 6394.41

Total cost per unit sold = 6396.41/390000 = 0.016 per Kg

d)

If Q = 15000 kg

Cost = (1-0.05)*6 = 5.7

Holding cost H = 0.25*5.7 = 1.425

Total Cost = Holding Cost + Ordering Cost = (Q/2)H + (D/Q)S

Total cost = (15000/2)*1.425 + (390000/15000)56 = 12143.5